OpenStudy (anonymous):

Express the following in roster form: Set N is the set of natural numbers between fourteen and nineteen.

OpenStudy (anonymous):

\[N\left\{ n|14\le n \le 19\right\} \]

OpenStudy (anonymous):

\[n \in N\]

OpenStudy (anonymous):

roster form is just a list {15,16,17,18} (as opposed to set builder {n in N | 14<n <19})